### v1

# Sharp Self-Normalized Concentration Inequalities of Marginal Mean with Sample Variance Only

Abstract
(This is the first version of a working paper. A more detailed follow-up with applications and uniform results is in preparation.) We develop a family of self-normalized concentration inequalities for marginal mean under martingale-difference structure and $\phi/\tilde{\phi}-$mixing conditions, where the latter includes many processes that are not strongly mixing. The variance term is fully data-observable: naive sample variance in the martingale case and an empirical block long-run variance under mixing conditions. Thus, no predictable variance proxy is required. No specific assumption on the decay of the mixing coefficients (e.g. summability) is needed for the validity. The constants are explicit and the bounds are ready to use.
1 Literature Review
This section is dedicated to the literature review of the previous research on self-normalized inequality and empirical Bernstein inequality for marginal mean of real-valued random variables.
1.1 Self-Normalized Inequality
Suppose $(\Omega,\mathcal{F},(\mathcal{A}_{i})_{i\in\mathbb{Z}},\mathbb{P})$ is a filter space and $(Z_{i})_{i\in\mathbb{Z}}=:\mathbf{Z}$ is a $\mathcal{Z}$-valued stochastic process adapted to $(\mathcal{A}_{i})_{i\in\mathbb{Z}}$, where $\mathcal{Z}$ is some Borel space. Denote
$\mathbf{M}_{n}=\sum_{i=1}^{n}(Z_{i}-\mathbb{E}[Z_{i}]),\ [\mathbf{M}]_{n}=\sum_{i=1}^{n}(Z_{i}-\mathbb{E}[Z_{i}])^{2},\langle\mathbf{M}\rangle_{n}=\sum_{i=1}^{n}\mathbb{E}[(Z_{i}-\mathbb{E}[Z_{i}])^{2}|\mathcal{A}_{i}].$
Generally speaking, the central goal of scalar self-normalized inequality is to show some concentration phenomenon of mean based on $\mathbf{M}_{n}$, $\mathbf{M}]_{n}$ and $\langle\mathbf{M}\rangle_{n}$. When $(Z_{i},\mathcal{A}_{i})_{i\in\mathbb{Z}}$ is a martingale difference sequence (MDS) such that $|Z_{i}|\leq b$, ? (?) did a pioneering work on this topic by showing an exponential upper bound of $\mathbb{P}(\mathbf{M}_{n}\geq x;\langle\mathbf{M}\rangle_{n}\leq v)$ for any $x,v\geq 0$. More specifically, when $\mathbb{E}[Z_{i}^{2}|\mathcal{A}_{i}]=\sigma^{2}>0$ holds, we have corollary
$\mathbb{P}\left(\frac{1}{n}\mathbf{M}_{n}\leq\sqrt{\frac{2\sigma^{2}\log(\frac{1}{\delta})}{n}}+\frac{b\log(\frac{1}{\delta})}{3}\right)\geq 1-\delta.$ (1)
Then, ? (?) and ? (?) extended Freedman’s work to unbounded cases. These results provide powerful concentration and limit theorems for Studentized partial sums and self-normalized martingales. However, the variance term typically appears as a predictable variance process or as an abstract conditional variance. which is rarely observable and often lives only as a process adapted to a given filtration. Recall that, when $(Z_{i},\mathcal{A}_{i})_{i\in\mathbb{Z}}$ is MDS, $\mathbb{E}[Z_{i}]=0$ holds for every $i$, which implies that $\mathbf{M}]_{n}=\sum_{i=1}^{n}Z_{i}^{2}$ is fully observable. Hence, deriving a concentration inequality purely based on $\mathbf{M}_{n}$ and $\mathbf{M}]_{n}$ becomes a meaningful task. By strengthening the assumption of MDS to symmetric MDS, ? (?) proved that, for all $x>0$,
$\mathbb{P}\Big(\frac{\mathbf{M}_{n}}{[\mathbf{M}]_{n}}\geq x\Big)\leq\sqrt{\mathbb{E}\Big[\exp(-\frac{1}{2}[\mathbf{M}]_{n}x^{2})\Big]}$ (2)
and, for all $x,y\geq 0$,
$\mathbb{P}\Big(\frac{\mathbf{M}_{n}}{[\mathbf{M}]_{n}}\geq x,[\mathbf{M}]_{n}\leq y\Big)\leq\exp\Big(-\frac{1}{2}x^{2}y\Big).$ (3)
Since conditionally symmetric MDS satisfies the definition of “heavy-on-left” in ? (?), they obtained the following important extension ?’s (2),
$\mathbb{P}\Big(\frac{\mathbf{M}_{n}}{[\mathbf{M}]_{n}}\geq x\Big)\leq\inf_{p\geq 1}\Big(\mathbb{E}\Big[\exp(-\frac{1}{2}(p-1)[\mathbf{M}]_{n}x^{2})\Big]\Big)^{\frac{1}{p}}.$ (4)
They also proved that (3) holds for “heavy-on-left” MDS as well. Compared with (1), (2)-(4) hold under much weaker moment conditions but their statistical applications are not so convenient as (1). the dependence on $\mathbf{M}]_{n}$ appears inside the Laplace transform or in the event $\{[\mathbf{M}]_{n}\leq y\}$, and there is no closed-form expression of the type $\sqrt{[\mathbf{M}]_{n}\log(\frac{1}{\delta})}$ with an explicit confidence parameter $\delta$. However, the variance ($\sigma$) and boundedness ($b$) are usually unknown to us, which reject the direct application of (1).
Under some circumstances, the value of boundedness can still be known from prior information. But, compared with boundedness, it is more challenging and risky to assume the variance to be known. Thus, it would be very ideal to derive an inequality having form (1) but the $\sigma$ is observable (or data-dependent), which is actually the core task of a parallel but tightly connected area, empirical Bernstein inequality of marginal mean.
1.2 Empirical Bernstein Inequality of Marginal Mean
Based on a variance estimator, ? (?) derived the following EB inequality for IID random variables taking values in $[0,1]$. However, compared with (1), their inequality is not sharp. i.e. The length of one sided confidence interval, denoted as $L_{n}$, does NOT satisfies
$\sqrt{n}L_{n}\xrightarrow[]{\mathbb{P}}\sqrt{2\log(1/\alpha)\sigma^{2}},$ (5)
where $\sigma^{2}$ is the marginal variance of these IID random variables. Based on self-normalized martingale techniques, Theorem 4 of ? (?) did a groundbreaking work by showing an anytime EB bound of the difference between sample mean and averaged conditional mean. i.e. Based on the $\mathbf{Z}$ introduced in Section 1.1, the following inequality holds for any given $\delta>0$,
$\mathbb{P}\Big(\forall\ n\geq 1:\Big|\frac{1}{n}\sum_{i=1}^{n}Z_{i}-\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[Z_{i}|Z_{i-1},...,Z_{1}]\Big|\leq\frac{u_{\delta}(V_{n})}{n}\Big)\geq 1-2\delta,$ (6)
where $(V_{i})_{i\geq 1}$ is a variance process adapted to filter $\{\sigma(Z_{1},...,Z_{i})\}_{i\geq 1}$. Function $u$ is essentially the sub-exponential uniform boundary and is user-defined (see Sections 2 and 3 of ? (?) for more details). It is obvious that, when $\mathbb{E}[Z_{i}|Z_{i-1},...,Z_{1}]$ is invariant with respect to $i$ (e.g. IID), (6) becomes an anytime EB inequality of marginal mean. But it has to pay for a iterated logarithmic loss of the sharpness since it is dedicated to be stopping-time valid. Thus, it is still open whether we can get a sharp EB inequality of marginal mean even for IID data. More recently, ? (?) (Theorem 2, Remark 1) gives a nice answer to this question (Their results include more general weighted empirical mean but we here only focus on simple empirical mean for the ease of comparison). By assuming $Z_{i}$’s as bounded random variables taking value in $[0,1]$, they proved the following EB inequality of marginal mean for all $\delta>0$ when $\mathbb{E}[Z_{i}|Z_{i-1},...,Z_{1}]=\mu$,
$\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}Z_{i}-\mu\leq\sqrt{\frac{2\log(\frac{1}{\delta})V_{n,\delta}}{n}}\right)\geq 1-\delta.$ (7)
Here $V_{n,\delta}$ is a quantity depending on sample and level $\delta$. More specifically, when the conditional variance is invariant, i.e., $\text{Var}(Z_{i}|Z_{i-1},...,Z_{1})=\sigma^{2}$, $V_{n,\delta}$ converges to $\sigma^{2}$ almost surely. Thus, (7) is sharp in the sense of (5). Up to our knowledge, ? (?) were the first ones to prove that their EB inequality is asymptotically sharp.
In summary, the results of Ramdas and co-authors are indeed elegant. But, as we discussed for fixed-n EB inequality of marginal mean, they have three structural limitations: (i) they require a known and tight bound on the support length to calculate the confidence interval; (ii) they primarily target the predictable mean (the arithmetic average of conditional expectations), and in general mixing settings there is no clear high-probability control of the gap between this predictable mean and the marginal mean—unless one assumes that all conditional expectations are constant, which essentially reduces to the martingale-difference case. Their recent important work, e.g. ? (?) and ? (?), still rely on the assumption of constant conditional expectation. (iii) they rely on a carefully tuned variance proxy satisfying the following three points simultaneously, (i) predictable; (ii) ensure the super-martingale property; (iii) converge to the true variance. The design of this proxy is highly non-trivial even for martingale differences or i.i.d. data. For some complex mixing processes, whose relevant variance of partial sum is the long-run variance rather than the marginal variance, the difficulty becomes even more significant. ? (?) derived EB inequalities for $\beta$-mixing Hilbert-valued processes based on block techniques. Their bounds depend explicitly on the knowledge of mixing coefficients and the block length, and the effective sample size is reduced to the number of blocks, which in general prevents parametric (see Theorems 1-3 there).
1.3 Our Contributions
We develop simple self-normalized (empirical Bernstein) inequalities for both martingale difference and mixing dependence structures. The main contributions of our results are:
  • C1
    (weak dependence on support information) Our bounds are of the form
    $|\text{Empirical Mean}-\text{Marginal Mean}|\lesssim\sqrt{\frac{2\log(\frac{1}{\delta})\text{Var}_{n}}{n^{2}}}\lor\frac{2bt}{n},$
    where $\text{Var}_{n}$ is some “empirical variance” and $b$ is the uniform bound of random variables. xcept in regimes of extremely small variance or very small sample size, the leading term is governed entirely by the empirical variance. In particular, in many practical situations one does not need a precise numerical value of the support length $b$; the bound automatically ignores $b$ once $\text{Var}_{n}$ is in a typical range.
  • C2
    (No variance-proxy selection: empirical variance only) Our self-normalization avoids the variance–proxy selection problem that is central in existing empirical Bernstein and betting martingale approaches. The bound depends only on a simple empirical variance: for martingale difference sequences this is the usual sample variance, and for mixing sequences it is a block-based empirical long-run variance. We do not require any predictable variance proxy, nor any tuning of a variance process to ensure a super-martingale property.
  • C3
    (Concentration for the marginal mean under complex dependence) We focus on concentration for the marginal mean in the presence of temporal dependence, rather than on the predictable (drifting) mean. In martingale-difference or i.i.d. settings these notions coincide, but in general mixing models, the gap between the two different kinds of means is hard to control. Our inequalities therefore are complement of the whole aforementioned important research of empirical Bernstein inequalities.
  • C4
    (No rate assumptions of mixing coefficients) Our inequalities are valid without any explicit assumptions of the decay rate of mixing coefficients, like summability. At the same time, the self-normalizer adapts to the relevant variance scale: in the martingale (or i.i.d.) case it corresponds to the marginal variance, whereas in the mixing case it estimates the long-run variance via a simple block construction.
  • C5
    (Asymptotically Sharp) Under very mild conditions, the $\text{Var}_{n}$ introduced in C1 consistently estimates the relevant variance parameter, $\sigma^{*}$. i.e. The marginal variance in the martingale-difference/i.i.d. case and the long-run variance in the mixing case. As a consequence, for any $\delta>0$,
    $\sqrt{n}\left(\sqrt{\frac{2\log(\frac{1}{\delta})\text{Var}_{n}}{n^{2}}}\lor\frac{2bt}{n}\right)\xrightarrow[]{\mathbb{P}}\sigma^{*}\sqrt{2\log(\frac{1}{\delta})}.$
    Therefore, when the assumption of constant conditional expectation holds, our inequalities are asymptotically sharp in the sense of (5).
2 Concentration Inequalities With Sample Variance Only
This section aims to exhibit our inequalities. Section 2.1 focus on martingale difference sequence and Section 2.2 focus on $\phi/\tilde{\phi}-$mixing processes.
2.1 Martingale Difference
In this subsection we first develop our self-normalized inequalities in the martingale-difference setting. As discussed in Section 1, most existing self-normalized concentration inequalities for the marginal mean either rely on the rarely observable predictable quadratic variation $\langle\mathbf{M}\rangle_{n}$,or are formulated for ratio-type statistics such as $\mathbf{M}_{n}/[\mathbf{M}]_{n}$. To the best of our knowledge, the following Theorem 1 provides the first sharp (in the sense of (5)) self-normalized concentration inequality for the marginal mean of a martingale difference sequence (MDS), with a fully observable variance term, which makes it also an empirical Bernstein inequality of MDS.
Theorem 1
Given a probability space $(\Omega,\mathcal{F},\mathbb{P})$, let $(\mathcal{A}_{i})_{i\geq 0}$ be a filter satisfying $\mathcal{A}_{i}\subset\mathcal{F}$. Suppose $(W_{i},\mathcal{A}_{i})_{i\geq 0}$ is a martingale difference sequence (MDS) such that $|Z_{i}|\leq b$ holds for any given $i\geq 1$. Then, by denoting $\mathbf{M}_{n}=\sum_{i=1}^{n}Z_{i}$, $\mathbf{M}]_{n}=\sum_{i=1}^{n}Z_{i}^{2}$, for any given $\delta>0$ and $n\geq 1$, we have
$\mathbb{P}\left(|\mathbf{M}_{n}|\leq\max\left\{\sqrt{2[\mathbf{M}]_{n}\log(\frac{1}{\delta})},2b\log(\frac{1}{\delta})\right\}\right)\geq 1-2\delta.$ (8)
As we can see in (8), combines the Bernstein–type tradeoff of Freedman’s inequality with a fully empirical self-normalization. It replaces the unobservable predictable quadratic variation $\langle\mathbf{M}\rangle_{n}$ by the quadratic variation $\mathbf{M}]_{n}$, which is completely determined by the data. Generally speaking, Theorem 1 will serve as the basic building block for our extensions to general mixing settings (Section 2.2), where $\mathbf{M}]_{n}$ is replaced by a block-based empirical long-run variance.
Of course, a more acceptable setting similar to MDS is that $\mathbb{E}[Z_{i}|\mathcal{A}_{i-1}]=\mu$ holds for every $i$, which makes $(Z_{i}-\mu,\mathcal{A}_{i})$ an MDS. However, when $\mu\neq 0$, $\mathbf{M}]_{n}$ becomes a term containing unknown parameter. Using empirical mean to replace $\mu$ yields the following empirical Bernstein inequality.
Corollary 1
Given a probability space $(\Omega,\mathcal{F},\mathbb{P})$, let $(\mathcal{A}_{i})_{i\geq 0}$ be a filter satisfying $\mathcal{A}_{i}\subset\mathcal{F}$. Suppose $(Z_{i})_{i\geq 0}$ is a process adapted to $(\mathcal{A}_{i})_{i\geq 0}$ and satisfies $\mathbb{E}[Z_{i}|\mathcal{A}_{i-1}]=\mu$ and $|Z_{i}|\leq b_{V}<\infty$. Then, by denoting $\overline{\mathbf{M}}_{n}=\frac{1}{n}\sum_{i=1}^{n}Z_{i}$ and $\mathbb{V}]_{n}=\sum_{i=1}^{n}(Z_{i}-\overline{\mathbf{M}}_{n})^{2}$, the following inequality holds for any given $t>0$ and $n\geq 1$,
$\mathbb{P}\Big(\Big|\overline{\mathbf{M}}_{n}-\mu\Big|\leq\nu_{n}(\delta)\sqrt{\frac{2[\mathbb{V}]_{n}\log(\frac{1}{\delta})}{n^{2}}}\lor\frac{2b_{V}\log(\frac{1}{\delta})}{n}\Big)\geq 1-2\delta,\ \ \nu_{n}(\delta)=\frac{1}{1-\sqrt{2\log(\frac{1}{\delta})/n}}.$ (9)
It is instructive to compare Corollary 1 with Theorem 2 and Remark 1 of ? (?). In the bounded i.i.d. (or martingale-difference) case, their empirical Bernstein confidence sequence is asymptotically sharp in the sense of (5): for fixed $\delta$, the $\sqrt{n}$-scaled radius converges to $\sigma\sqrt{2\log(\frac{1}{\delta})}$. They work in a time-uniform setting and therefore construct a predictable variance proxy based on one-step-ahead predictions, together with a nontrivial betting/mixture strategy, and their target is the predictable (drifting) mean, which coincides with the marginal mean only under the martingale-difference assumption. Moreover, the sharpness discussion in Remark 1 is formulated under a homoscedasticity condition, where the conditional variance process converges to a constant so that the variance parameter is a single $\sigma^{2}$. The heteroscedastic case, with $\text{Var}(Z_{i})$ varying in $i$, is not treated explicitly. By contrast, Corollary 1 is purely fixed-$n$, focuses directly on the marginal mean, and uses only the naive empirical variance (plus a ”soft floor” depending on the range $b$, with no predictable variance proxy, no betting construction, and no tuning. For bounded independent or martingale-difference sequences with possibly non-constant marginal variances, the empirical variance automatically converges to the natural average scale $\frac{1}{n}\sum_{i=1}^{n}\text{Var}(Z_{i})=:\sigma^{2}_{n}$, and simultaneously, the $\sqrt{n}$-scaled radius converges to $\sigma_{n}\sqrt{2\log(\frac{1}{\delta})}$. In this sense, our empirical Bernstein bound is asymptotically sharp for general bounded independent or martingale-difference sequences, not only in the homoscedastic setting considered in ?’s sharpness discussion. Meanwhile, except for some extreme cases like significantly small variance and sample size, we can ignore the boundedness $b$ in high-probability bound, which is quite useful in the cases where we do not know the exact value of boundedness.
However, many practitioners may be inclined to ignore the $\frac{2}{n}b_{V}\log(\frac{1}{\delta})$ directly. A natural question is the cost of this ignorance. Corollary 2 shows that, in a non-asymptotic level, the ignorance of $\frac{2}{n}b_{V}\log(\frac{1}{\delta})$ only cost an exponentially small loss of coverage.
Corollary 2
Suppose $(Z_{i})_{i=1}^{n}$ are independent copies of non-degenerated random variable $Z_{0}$ such that $|Z_{0}|\leq b$. By denoting $\text{Var}(Z_{0})=:\sigma^{2}$, $E|Z_{0}-\mu|^{4}=:m_{4}$,
$\mathbb{P}\left(\Big|\overline{\mathbf{M}}_{n}-\mu\Big|\leq\nu_{n}(\delta)\sqrt{\frac{2\log(\frac{1}{\delta})[\mathbb{V}]_{n}}{n^{2}}}\right)\geq 1-2\delta-\exp\left(-\frac{n(1-\eta)^{2}\sigma^{4}}{2(m_{4}+\frac{1}{3}b(1-\eta)\sigma^{2})}\right)$
holds for every $\delta>0$, $\eta\in(0,1)$ and $n\geq N(\delta):=\min\{n\geq 1:\eta\sigma^{2}\geq 2\frac{1}{n}\log(\frac{1}{\delta})b_{V}^{2}\}$, where $\nu_{n}(\delta)$ is introduced in Corollary 1.
To summarize, in the martingale-difference and i.i.d. regimes our inequalities provide fully data-driven, asymptotically sharp empirical Bernstein bounds for the marginal mean, with the variance term given by the usual sample variance. In the next subsection we move beyond martingale differences and develop analogous self-normalized inequalities for $\phi/\tilde{\phi}-$mixing processes, where the relevant variance scale becomes the long-run variance and is estimated by a simple block-based empirical long-run variance.
2.2 $\phi$ and $\tilde{\phi}-$Mixing
2.2.1 Review of $\phi/\tilde{\phi}-$mixing Conditions
Based on the process $\mathbf{Z}$ introduced in Section 1.1, $\phi$-mixing condition is defined as follow.
Definition 1
By letting $\mathcal{F}_{a}^{b}$ be the sigma algebra generated by $Z_{i}$’s whose $i\in[a,b]$, we say process $\mathbf{Z}$ satisfies $\phi$-mixing condition (or is a $\phi$-mixing process) if
$\phi(k):=\max_{n}\sup_{A\in\mathcal{F}_{-\infty}^{n},\mathbb{P}(A)>0}\sup_{B\in\mathcal{F}_{k+n}^{\infty}}|\mathbb{P}(B|A)-\mathbb{P}(B)|$ (10)
converges to 0 as $k\nearrow\infty$. We address $\phi(k)$ as the $\phi$-mixing coefficient.
As pointed out by ? (?), under some proper Doeblin conditions, $\phi$-mixing condition (see (1)) includes many stationary Markov chains. A good example is AR(1) model with standard Gaussian white noise and the absolute value of auto-regression coefficient smaller than 1. For more thorough discussion, please refer to textbook ? (?). It is widely known that $\phi$-mixing coefficient is more restrictive than $\alpha$-mixing condition (see ? (?)), since $\phi(k)\geq\alpha(k)$ holds for any $k\geq 0$. However, the generality of $\alpha$-mixing coefficients is still not strong enough to cover many fundamental real-valued dynamic systems and time series models, since their $\alpha$-mixing coefficients do not converge to 0. A simple but non-trivial example is $X_{t}=\sum_{j=1}^{\infty}2^{-j}\varepsilon_{t-j}$, where $\{\varepsilon_{t}\}_{t\in\mathbb{Z}}$ is independent Bernoulli process with parameter $\frac{1}{2}$. Obviously, this is the stationary solution of AR(1) model $X_{t}=\frac{1}{2}X_{t-1}+\frac{1}{2}\varepsilon_{t}$. ? (?) shew that the $\alpha$-mixing coefficient of this process is larger or equal to $\frac{1}{4}$. Regarding $\phi$-mixing condition is a special case of $\alpha$-mixing condition, $\phi$-mixing condition also fails to incorporate this fundamental first-order auto-regression model.
? (?) discovered that this real-valued AR(1) process satisfies the following mixing condition. (? still named this new mixing condition as $\phi$ in their paper, while in their later textbook, ? (?), they name it as $\tilde{\phi}$. To distinguish the difference, we here chose the later name.)
Definition 2
Suppose $\mathcal{Z}\subset\mathbb{R}$. We say $\mathbf{Z}$ satisfies $\tilde{\phi}-$mixing condition (or is a $\tilde{\phi}-$mixing process) if
$\tilde{\phi}(k):=\sup_{z\in\mathbb{R}}||\mathbb{P}(Z_{n+k}\leq z|\mathcal{F}_{-\infty}^{n})-\mathbb{P}(Z_{n+k}\leq z)||_{\infty}$ (11)
converges to 0 as $k\nearrow\infty$. We address $\tilde{\phi}(k)$ as the $\tilde{\phi}-$mixing coefficient.
Expect for the previous specific example, there are many other important real-valued time series models satisfying $\tilde{\phi}-$mixing conditions, like nonlinear causal shifts with independent innovations and smooth marginal distributions, non-linear and non-separable auto-regressive models with Lipschitz continuous recurrent function. Actually, under mild conditions, some interesting dynamic systems satisfying $\tilde{\phi}-$mixing condition as well. For more specific knowledge, please refer to Chapters 2 and 3 of ? (?).
2.2.2 Inequalities
For each given $n\geq 1$, let $\mathcal{H}_{n}:=\{h_{n}:\mathcal{Z}\to\mathbb{R}\}$ is a class of measurable mappings such that $h_{n}\in[a_{n},b_{n}]\subset\mathbb{R}$ holds for any given $n$ and $\mathbb{E}[h_{n}(Z_{i})]=\mu_{n}$ holds for any given $i$ and $n$. As for the dependence, we majorly consider the cases where process $\mathbf{Z}$ satisfies $\phi$ or $\tilde{\phi}-$mixing conditions introduced as follow. By letting $\mathbb{P}_{n}(h_{n})=\frac{1}{n}\sum_{i=1}^{n}h_{n}(Z_{i})$, this main goal of section is to give self normalized concentration inequality of $|\mathbb{P}_{n}(h_{n})-\mu|$ for $\phi/\tilde{\phi}-$mixing processes without assuming summability.
Theorem 2
($\phi$-mixing) Suppose $\mathbf{Z}$ is a $\mathcal{Z}$-valued $\phi$-mixing process and denote $\Phi_{n}=\sum_{k=1}^{n}\phi(k)$. For any given $l>0$ such that $\lim_{n}l\nearrow\infty$ and $l=o(\sqrt{n})$, by letting $m=[\frac{n}{[l]}]$, we have partition $\{1,...,n\}=\cup_{j=1}^{m+1}B_{j}$. We also define $\bar{H}_{n}=\frac{1}{m[l]}\sum_{i=1}^{m[l]}h_{n}(Z_{i})$. Then, for arbitrary $\xi_{n}\searrow 0$ and any given $\delta>0$, $n\geq 1$, we have
$\mathbb{P}\left(\Big|\frac{1}{n}S_{nn}(h_{n})\Big|\leq\tilde{\nu}_{n}(\delta)\max\{A_{\phi},B_{\phi}\}+\frac{[l]}{n}(b_{n}-a_{n})\right)\geq 1-2\delta$
where
$\tilde{\nu}_{n}(\delta)=\Big(1-\sqrt{\frac{2\log(\frac{1}{\delta})}{m}}\Big)^{-1},B_{\phi}=\frac{2([l]+2\Phi_{n})(b_{n}-a_{n})}{n},$
$A_{\phi}$ $=\sqrt{2\log(\frac{1}{\delta})\frac{1}{n}\hat{V}_{n}}+4(b_{n}-a_{n})\Phi_{n}\sqrt{\frac{2\log(\frac{1}{\delta})m}{n^{2}}}+\sqrt{\frac{2\log(\frac{1}{\delta})\xi_{n}}{n^{2}}},$
$\hat{V}_{n}=\frac{1}{n}\sum_{j=1}^{m}(\sum_{i\in B_{i}}(h_{n}(Z_{ni})-\bar{H}_{n}))^{2}.$
Theorem 3
($\tilde{\phi}-$mixing) Suppose $\mathbf{Z}$ is a $\mathbb{R}$-valued $\tilde{\phi}-$mixing process and denote $\tilde{\Phi}_{n}=\sum_{k=1}^{n}\tilde{\phi}(k)$. We further assume that $h_{n}||_{TV}<\infty$ holds for each given $n$, where $||\cdot||_{TV}$ is the total variation norm of $h_{n}:\mathcal{Z}\to\mathbb{R}$. Then, for arbitrary $\xi_{n}\searrow 0$ and any given $\delta>0$, $n\geq 1$,
$\mathbb{P}\left(\Big|\frac{1}{n}S_{nn}(h_{n})\Big|\leq\tilde{\nu}_{n}(\delta)\max\{A_{\tilde{\phi}},B_{\tilde{\phi}}\}+\frac{[l]}{n}(b_{n}-a_{n})\right)\geq 1-2\delta,$
[…]
where $\tilde{\nu}_{n}(\delta)$ and $\hat{V}_{n}$ are introduced in Theorem 2.
Remark 1
A noteworthy point is that $l$ here is a user-defined parameter, which is totally different from the size of block in the widely used blocking skills (e.g. ? (?)). For arbitrary chosen $l\nearrow\infty$ such that $l=o(\sqrt{n})$, the sharpness of the high probability bounds exhibited in Theorems 2 and 3 do not change.
Theorems 2 and 3 extend the martingale-difference bound of Theorem 1 to weakly dependent time series. Under mild $\phi/\tilde{\phi}-$mixing assumptions, we obtain self-normalized Bernstein-type inequalities in which the variance term is a block-based empirical long-run variance $\hat{V}_{n}$ that is fully observable from the data. The structure of the bounds is exactly analogous to the martingale case, in which the leading self-normalized term $\sqrt{\frac{2\hat{V}_{n}}{n}\log(\frac{1}{\delta})}$ and a coverage probability $1-\delta$. In particular, we do not impose or use any explicit decay rate for the mixing coefficients. Mixing coefficient affects the high-probability bound in a very weak way. Conceptually, these results provide fixed-$n$ empirical Bernstein inequalities for the marginal mean of weakly dependent sequences, with a variance scale that automatically adapts to the appropriate long-run variance whenever it exists. This contrasts with classical Bernstein-type inequalities for mixing processes, which typically involve non-empirical long-run variance bounds and explicit mixing coefficients, or pay a loss in rate due to blocking. Here the only variance term is the simple block-based empirical long-run variance, and the resulting bounds retain a parametric $\sqrt{n}$-rate and the same sharp constant $\sqrt{2\log(\frac{1}{\delta})}$ as in the martingale-difference case. The knowledge of mixing coefficient affects the construction high-probability bound in a very indirect way. For example, under many circumstances, it is natural to assume the mixing coefficients are summable, which means $\max_{n}\Phi_{n}\lor\tilde{\Phi}_{n}<\infty$. Thus, replacing them with any user-defined diverging sequence, like $\log\log n$, yield a valid high-probability bound and the sharpness is not harmed at all since leading term is still $\sqrt{\frac{2\hat{V}_{n}}{n}\log(\frac{1}{\delta})}$.
Unfortunately, we sometimes have to face some “strongly-dependent” situations where $\lim_{n}\sum_{k=1}^{n}\tilde{\phi}(k)<\infty$ (or $\lim_{n}\sum_{k=1}^{n}{\phi}(k)<\infty$) does NOT hold, like $\tilde{\phi}(k)=1/k$. Under this circumstance, we can not even guarantee the summability of autocovariances and the existence of the classical long-run variance. A more severe question is that it is often difficult to detect diverging speed of $\sum_{k=1}^{n}\tilde{\phi}(k)$ and $\sum_{k=1}^{n}\phi(k)$. Hence, similar to our discussion in Section 2.1, we can get out of this awkward situation by ignoring the terms containing unknown parameters. The following Theorem 4 precisely describe the cost of this ignorance. Similar to Corollary 2, the cost is usually on an exponential-level, which is often quite affordable.
Theorem 4
Suppose $\mathbf{Z}$ is a $\mathcal{Z}$-valued $\phi/\tilde{\phi}-$mixing process. Based on the notations used in Theorems 2 and 3, for all $\delta>0$, $n\geq 1$ and user-defined $\theta^{\prime}\in(0,\frac{1}{2})$, $c_{n}\searrow 0$ and $t_{n}\searrow 0$, we have
$\mathbb{P}\Big(|\frac{1}{n}S_{nn}(h_{n})|\leq\tilde{\nu}_{n}(\delta)\max\{\mathbf{A}^{\prime},\mathbf{B}^{\prime}\}+c_{n}\Big)\geq 1-2\delta-\sum_{k=1}^{3}\text{Error}_{k},$
where $\mathbf{A}^{\prime}=\sqrt{2\frac{1}{n}\log(\frac{1}{\delta})\hat{V}_{n}}+t_{n}\sqrt{2\log(\frac{1}{\delta})}+\sqrt{\frac{2\xi_{n}\log(\frac{1}{\delta})}{n^{2}}},$
$\mathbf{B}^{\prime}=\frac{2\log(\frac{1}{\delta})[l](b_{n}-a_{n})}{n}+\frac{\log(\frac{1}{\delta})}{n^{1-\theta^{\prime}}},\ \text{Error}_{1}=2\exp\left(-2\Big(\frac{(n/[l])c_{n}}{(b_{n}-a_{n})}\Big)^{2}\right),$
$\text{Error}_{2}=2m\exp\left(-0.5\Big(\frac{\sqrt{n[l]}t_{n}}{\triangle_{n}\Psi_{n}}\Big)^{2}\right),\ \text{Error}_{3}=2m\exp\left(-\Big(\frac{n^{\theta^{\prime}}}{16\triangle_{n}\Psi_{n}}\Big)^{2}\right).$
Here $\Psi_{n}$ indicates $\Phi_{n}$ and $\widetilde{\Phi}_{n}$ under $\phi$ and $\tilde{\phi}-$mixing conditions respectively. $\triangle_{n}$ means $2(b_{n}-a_{n})$ and $h_{n}||_{TV}$ under $\phi$ and $\tilde{\phi}-$mixing conditions respectively
Suppose $a_{n}=a$ and $b_{n}=b$. For $\phi$-mixing process with coefficient $\phi(k)=\frac{1}{k}$, the ignorance of parameters $\Phi_{n}$ is nearly harmless. Generally speaking, even in regimes where the autocovariances are not summable and a classical long-run variance may fail to exist, Theorem 4 still delivers stable, fully data-driven high-probability interval based on the block empirical variance. In such cases we may not have any CLT-type interpretation, but the bounds remain valid and operational, providing finite-sample control of the marginal mean under very strong dependence.
3 Conclusions
In this paper we have developed a family of self-normalized concentration inequalities for marginal means, in both martingale-difference and weakly dependent settings, with variance terms that are fully observable from the data. In the martingale-difference (and in particular i.i.d.) case, our bounds take the form of empirical Bernstein inequalities where the self-normalizer is the usual sample variance. In the $\phi/\tilde{\phi}-$mixing case, the self-normalizer becomes a simple block-based empirical long-run variance. In all cases, the bounds have an explicit Freedman/Bernstein-type structure: a leading term of order $\sqrt{\text{empirical (long run) variance}\log(\frac{1}{\delta})/n}$ and a soft floor of order $\frac{b}{n}\log(\frac{1}{\delta})$, with coverage probability $1-\delta$ up to an exponentially small remainder.
We view these results as a foundational step towards a broader theory of sharp, fully empirical self-normalized inequalities for dependent data. A more detailed follow-up, including uniform-in-time versions, statistical applications, and high-dimensional extensions, is in preparation.
Appendix A Proof of Main Results and Technical Lemmas
A.1 Technical Lemmas
Lemma A.1
Given any two positive integers $m$ and $l$, define $\{1,...,ml\}=\bigcup_{j=1}^{m}B_{j}$, where $B_{j}=\{(j-1)l+1,...,jl\}$. Suppose $\{Z_{i}\}_{i=1}^{ml}$ is a group of real numbers. By denoting $\bar{w}=\frac{1}{ml}\sum_{i=1}^{ml}Z_{i}$, we have the following identity holds for every $m,l\geq 1$,
$\sum_{j=1}^{m}(\sum_{i\in B_{j}}(Z_{i}-\bar{w}))^{2}+l^{2}m(\bar{w}-\mu)^{2}=\sum_{j=1}^{m}(\sum_{i\in B_{j}}(Z_{i}-\mu))^{2}$
Lemma A.2
Suppose $\mathcal{A}$ and $\mathcal{B}$ are two $\sigma$-algebras. By denoting $\mathcal{R}=\{A\cap B:A\in\mathcal{A},B\in\mathcal{B}\}$, we have
$\sigma(\mathcal{R})=\sigma(\mathcal{A}\cup\sigma(\Lambda))=:\mathcal{A}\lor\sigma(\Lambda),$
where $\sigma(\mathcal{R})$ and $\sigma(\Lambda)$ are the sigma algebras generated by set $\mathcal{R}$ and random variable $\Lambda$.
Lemma A.3
Suppose $(\Omega,\mathcal{F},\mathbb{P})$ is a probability space and $\mathcal{A}\subset\mathcal{F}$ is a sub $\sigma$-algebra. Suppose $\Lambda$ is a random variable taking value in Borel measurable space $(E,\mathcal{B}(E))$. Provided that $\phi:E\times\Omega\to[0,\infty)$ is measurable with respect to $\mathcal{B}(E)\otimes\mathcal{A}$,
$\mathbb{E}[\phi(\Lambda(\cdot),\cdot)|\mathcal{A}](\omega)=\int_{E}\phi(\lambda,\omega)\mu_{\omega}(d\lambda)\ \ \text{a.s.},\$(A.1)
where $\mu_{\omega}(\cdot)=\mathbb{P}(\Lambda\in\cdot\ |\mathcal{A})(\omega)$ is the regular conditional distribution of random variable $\Lambda$ on condition of $\sigma$-algebra $\mathcal{A}$, which, for any fixed $\omega$, is a measure defined on Borel field $\mathcal{B}(E)$.
Lemma A.4
Based on the conditions introduced in Lemma A.2, suppose mapping $h:E\times\Omega\to[0,\infty]$ is measurable with respect to $\mathcal{B}(E)\times\mathcal{F}$. Meanwhile, for every fixed $\lambda\in E$, by letting $g(\lambda,\omega):=E[h(\lambda,\cdot)|\mathcal{A}](\omega)$ and $\mathcal{G}=\mathcal{A}\lor\sigma(\Lambda)$, we have
$\mathbb{E}[h(\Lambda(\cdot),\cdot)|\mathcal{G}](\omega)=g(\Lambda(\omega),\omega)\ \ a.s..$
Remark A.2
$g(\lambda,\omega):=E[h(\lambda,\cdot)|\mathcal{A}](\omega)$ is random variable measurable with respect to $\mathcal{A}$ but, for any fixed $\omega$, it is a deterministic mapping of $\lambda$.
Lemma A.5
Suppose $\mathbf{Z}$ is a $\mathcal{Z}$-valued $\tilde{\phi}-$mixing process defined as Definition 1. Let $\mathcal{H}_{n}$ be a class of real-valued and measurable functions defined on $\mathcal{Z}$ and $h_{n}\in[a_{n},b_{n}]$. Then,
$||\mathbb{E}[h_{n}(Z_{n+k})|\mathcal{F}_{-\infty}^{n}]-\mathbb{E}[h_{n}(Z_{n+k})]||_{\infty}\leq 2(b_{n}-a_{n})\phi(k)$
for any given $n$ and $k$. Furthermore, suppose $\mathbf{Z}$ is a $\mathcal{Z}$-valued $\tilde{\phi}-$mixing process defined as Definition 2 and $\mathcal{Z}\subset\mathbb{R}$. Then, by denoting $h_{n}||_{TV}$ as the total-variation norm of $h_{n}$,
$||\mathbb{E}[h_{n}(Z_{n+k})|\mathcal{F}_{-\infty}^{n}]-\mathbb{E}[h_{n}(Z_{n+k})]||_{\infty}\leq||h_{n}||_{TV}\tilde{\phi}(k),\ \ \forall\ n,k\geq 1$
Lemma A.6
Suppose $(\Omega,\mathcal{F},(\mathcal{A}_{i})_{i\in\mathbb{Z}},\mathbb{P})$ is a filter space and $(W_{i})_{i\in\mathbb{Z}}$ is a $\mathcal{W}$-valued process adapted to $(\mathcal{A}_{i})_{i\in\mathbb{Z}}$. Suppose $f:\mathcal{W}\to\mathbb{R}$ is a measurable mapping such that $\mathbb{E}[f(W_{i})]=\mu_{f}$. Then, provided that the projective criteria $\max_{i\in\mathbb{Z}}\sum_{k>i}||\mathbb{E}[(f(W_{k})-\mu_{f})|\mathcal{A}_{i}]||_{p}<\infty$ holds for some $p\geq 1$, the following identity holds
$f(W_{i})$ $=D_{i}(f)-\varepsilon_{i}(f)+\varepsilon_{i-1}(f),$
$S_{n}(f):$ $=\sum_{i=1}^{n}f(W_{i})=\sum_{i=1}^{n}D_{i}(f)-\varepsilon_{n}(f)+\varepsilon_{0}(f),$
where $D_{i}(f)$, $\varepsilon_{i}(f)$ are mean-zero random variables with finite $L^{p}$ norm. Furthermore, we have
$D_{i}=\sum_{k\geq i}(\mathbb{E}[f(W_{k})|\mathcal{A}_{i}]-\mathbb{E}[f(W_{k})|\mathcal{A}_{i-1}]),\ \varepsilon_{i}=\sum_{k>i}\mathbb{E}[(f(W_{k})-\mu_{f})|\mathcal{A}_{i}].$
Proof of Lemmas A.1 and A.2 The results of Lemma A.1 are just natural consequence of some simple algebra. We thus omit it here. Lemma A.2 is a classical result widely discussed in many measure theory textbooks.
Proof of Lemmas A.3 Our method of proof is still the classical “testing function method”, which is widely used in measure theory.
(Step 0) (Being measurable and well-defined) Regarding $\phi$ is measurable with respect to $\mathcal{B}(E)\otimes\mathcal{A}$, for any given $\lambda\in E$, $\phi(\lambda,\omega)$ is measurable with respect to $\mathcal{A}$. Thus, the right hand side of (A.1) is a well defined random variable.
(Step 1) (Simple Function) Suppose $\{A_{i}\}_{i=1}^{n}\subset\mathcal{A}$ and $\{E_{j}\}_{j=1}^{n}\subset\mathcal{B}(E)$ are partitions of $\Omega$ and $E$. Suppose $\phi$ is a simple function of form
$\phi_{n}(\lambda,\cdot):=\sum_{i=1}^{n}c_{i}1_{E_{i}\times A_{i}}(\lambda,\cdot)=\sum_{i=1}^{n}c_{i}1_{E_{i}}(\lambda)1_{A_{i}}(\cdot).$ (A.2)
Then, the following identities hold at least almost surely,
$\mathbb{E}[\phi_{n}(\Lambda(\cdot),\cdot)|\mathcal{A}](\omega)=\sum_{i=1}^{n}c_{i}\mathbb{E}[1_{E_{i}}(\Lambda(\cdot))1_{A_{i}}(\cdot)|\mathcal{A}](\omega)$
$=$ $\sum_{i=1}^{n}c_{i}1_{A_{i}}(\omega)\mathbb{E}[1_{E_{i}}(\Lambda(\cdot))|\mathcal{A}](\omega)=\sum_{i=1}^{n}c_{i}1_{A_{i}}(\omega)\int_{E}1_{E_{i}}(\lambda)\mu_{\omega}(d\lambda)$
$=$ $\int_{E}\sum_{i=1}^{n}c_{i}1_{A_{i}}(\omega)1_{E_{i}}(\lambda)\mu_{\omega}(d\lambda)=\int_{E}\phi_{n}(\lambda,\omega)\mu_{\omega}(d\lambda).$ (A.3)
(Step 2) (Non-negative Function) It is widely known that, for any given measurable $\phi$, there exists a sequence of simple functions, denoted $\{\phi_{n}\}$, such that $\lim_{n\to\infty}\phi_{n}\nearrow\phi$ pointwisely and each $\phi_{n}$ is of form (A.2). Then, by using conditional monotonous convergence theorem and (A.1), we obtain
$\mathbb{E}[\phi(\Lambda(\cdot),\cdot)|\mathcal{A}](\omega)=\lim_{n\to\infty}\mathbb{E}[\phi_{n}(\Lambda(\cdot),\cdot)|\mathcal{A}]=\lim_{n\to\infty}\int_{E}\phi_{n}(\lambda,\omega)\mu_{\omega}(d\lambda)=\int_{E}\phi(\lambda,\omega)\mu_{\omega}(d\lambda).$
Q.E.D
Proof of Lemma A.4 Similar to Lemma A.3, the proof is also divided into multiple steps.
Step 1 (Extending the measurability) Regarding that we do not know whether $g$ is measurable with respect to $\mathcal{B}(E)\otimes\mathcal{A}$, in this step, we focus on showing the existence of a mapping $(\lambda,\omega)\to\tilde{g}(\lambda,\omega)$ such that (i) $\tilde{g}=g$ holds almost surely; (ii) $(\lambda,\omega)\to\tilde{g}(\lambda,\omega)$ is measurable with respect to $\mathcal{B}(E)\otimes\mathcal{A}$.
Note that, for any given $\omega\in\Omega$ and partitions $\{E_{k}\}_{k=1}^{K}\subset\mathcal{B}(E)$, $\{F_{k}\}_{k=1}^{K}\subset\mathcal{A}$ , by denoting $h_{K}(\lambda,\cdot)=\sum_{k=1}^{K}a_{k}1_{E_{k}}(\lambda)1_{F_{k}}(\cdot)$, $a_{k}\geq 0$, we have
$\mathbb{E}[h_{K}(\lambda,\cdot)|\mathcal{A}](\omega)=\sum_{k=1}^{K}a_{k}\mathbb{E}[1_{E_{k}}(\lambda)1_{F_{k}}(\cdot)|\mathcal{A}](\omega)=\sum_{k=1}^{K}a_{k}1_{E_{k}}(\lambda)\mathbb{E}[1_{F_{k}}(\cdot)|\mathcal{A}](\omega)=:\tilde{g}_{K}(\lambda,\omega).$
$\tilde{g}_{K}(\lambda,\omega)$ is obviously measurable with respect to $\mathcal{B}(E)\times\mathcal{A}$. Then, for any non-negative mapping $h$ measurable with respect to $\mathcal{B}(E)\otimes\mathcal{A}$, there exists a sequence of non-negative simple functions $\{h_{K}\}$ such that (i) $h_{K}\leq h_{K+1}$; (ii) $\lim_{K\to\infty}h_{K}=h$ pointwisely. Then, for any given $\lambda\in E$ and $\omega\in\Omega$,
$\mathbb{E}[h(\lambda,\cdot)|\mathcal{A}](\omega)=\lim_{K\to\infty}\mathbb{E}[h_{K}(\lambda,\cdot)|\mathcal{A}](\omega)=\lim_{K\to\infty}g_{K}(\lambda,\omega)=:\tilde{g}(\lambda,\omega).$ (A.4)
Then, $\tilde{g}(\lambda,\omega)$ is measurable with respect to $\mathcal{B}(E)\otimes\mathcal{A}$ and $\tilde{g}(\lambda,\omega)=g(\lambda,\omega)$ holds almost surely according to the uniqueness of conditional expectation.  [^1] Step 1 is actually dedicated to show that we can use conditional Fubini theorem with kernel (see Step 2) without worrying about measurability of $g$.
Step 2 Define $\mathcal{R}=\{A\cap\{\Lambda\in B\}:A\in\mathcal{A},B\in\mathcal{B}(E)\}$ and note that Lemma A.2 implies
$\sigma(\mathcal{R})=\sigma(\Lambda)\lor\mathcal{A}=:\mathcal{G}.$ (A.5)
In order to avoid unnecessary confusion of notations, for any real-valued random variable $Z:\Omega\to\mathbb{R}$, we also use $\mathbb{E}[Z(\omega)]:=\int_{\Omega}Z(\omega)d\mathbb{P}$ to indicate its expectation. Unlike conventional notation of expectation, our notation highlights the fact that a real-valued random variable is also a measurable function whose domain is $\Omega$, which could avoid many unnecessary confusion of notations later. More specifically, the main goal of this step is to prove that
$\mathbb{E}[1_{C}(\omega)\tilde{g}(\Lambda(\omega),\omega)]=\mathbb{E}[1_{C}(\omega)h(\Lambda(\omega),\omega)],\ \ \forall\ C\in\mathcal{R}.$ (A.6)
First, the definition of $\mathcal{R}$ yields
$\mathbb{E}[1_{C}(\omega)\tilde{g}(\Lambda(\omega),\omega)]=\mathbb{E}[1_{A}(\omega)1_{B}(\Lambda(\omega))\tilde{g}(\Lambda(\omega),\omega)]=\mathbb{E}[1_{A}(\omega)E[1_{B}(\Lambda(\cdot))\tilde{g}(\Lambda(\cdot),\cdot)|\mathcal{A}](\omega)].$
Then, by considering $1_{B}(\Lambda(\cdot))\tilde{g}(\Lambda(\cdot),\cdot)$ as the $\phi(\Lambda(\cdot),\cdot)$” in Lemma A.3, we have the following identities,
$\mathbb{E}[1_{A}(\omega)E[1_{B}(\Lambda(\cdot))\tilde{g}(\Lambda(\cdot),\cdot)|\mathcal{A}](\omega)]=\mathbb{E}[1_{A}(\omega)\int_{B}\tilde{g}(\lambda,\omega)\mu_{\omega}(d\lambda)]$
$=$ $\mathbb{E}[1_{A}(\omega)\int_{B}\mathbb{E}[h(\lambda,\cdot)|\mathcal{A}](\omega)\mu_{\omega}(d\lambda)]=\mathbb{E}[1_{A}(\omega)\mathbb{E}[\int_{B}h(\lambda,\cdot)\mu_{\omega}(d\lambda)|\mathcal{A}](\omega)],$
where the first to the last equations are because of Lemma A.3, Step 1 and “conditional Fubini theorem with kernel” respectively. Together with law of iterated expectation, we finish the proof of equation (A.6).
Step 3 Recall that $\mathcal{G}=\mathcal{B}(E)\otimes\mathcal{A}$. By defining
$\mathcal{H}=\{G\in\mathcal{G}:\mathbb{E}[1_{G}(\omega)\tilde{g}(\Lambda(\omega),\omega)]=\mathbb{E}[1_{G}(\omega)h(\Lambda(\omega),\omega)]\},$
results of Step 2 asserts $\mathcal{R}\subset\mathcal{H}\subset\mathcal{G}=\sigma(\mathcal{R})$. Now we aim to show $\mathcal{H}$ is a Dynkin system. i.e. (1) $\Omega\in\mathcal{H}$; (2) If $H_{1}\subset H_{2}\in\mathcal{H}$, $H_{2}\backslash H_{1}\in\mathcal{H}$; (3) If $\{H_{n}\}_{n\geq 1}$ is an non-decreasing sequence in $\mathcal{H}$, $\bigcup_{n\geq 1}H_{n}\in\mathcal{H}$. Since (1) and (2) are obvious, we only specify (3).
To prove (3), we first note that $\lim_{n\to\infty}1_{H_{n}}(\omega)=1_{H^{*}}(\omega)$ holds pointwisely on $\Omega$, where $H^{*}=\bigcup_{n\geq 1}H_{n}$. Then, using monotonous convergence theorem yields
$\mathbb{E}[1_{H^{*}}(\omega)\tilde{g}(\Lambda(\omega),\omega)]=\lim_{n\to\infty}\mathbb{E}[1_{H_{n}}(\omega)\tilde{g}(\Lambda(\omega),\omega)]=\lim_{n\to\infty}\mathbb{E}[1_{H_{n}}(\omega)h(\Lambda(\omega),\omega)]=\mathbb{E}[1_{H^{*}}(\omega)h(\Lambda(\omega),\omega)].$
Thus, (3) holds and we manage to show $\mathcal{H}$ is a Dynkin system. Meanwhile, since $\mathcal{R}$ is obviously a $\pi$-system, together with Sierpiński–Dynkin’s Theorem, we have
$\mathcal{G}:=\sigma(\mathcal{R})=\text{Dy}(\mathcal{R})\subset\mathcal{H}\subset\sigma(\mathcal{R})=:\mathcal{G},$
where $\text{Dy}(\mathcal{R})$ is the Dynkin system generated by $\pi$-system $\mathcal{R}$. (Some textbooks of measure theory also address Dynkin system as “Lambda” system.) Consequently, this implies that, for every $G\in\mathcal{G}$,
$\mathbb{E}[1_{G}(\omega)\tilde{g}(\Lambda(\omega),\omega)]=\mathbb{E}[1_{G}(\omega)h(\Lambda(\omega),\omega)].$ (A.7)
Step 4 Notice that, for any $G\in\mathcal{G}:=\sigma(\Lambda)\lor\mathcal{A}$,
$\mathbb{E}[1_{G}(\omega)g(\Lambda(\omega),\omega)]=\mathbb{E}[1_{G}(\omega)\tilde{g}(\Lambda(\omega),\omega)]=\mathbb{E}[1_{G}(\omega)h(\Lambda(\omega),\omega)]=\mathbb{E}[1_{G}(\omega)\mathbb{E}[h(\Lambda(\cdot),\cdot)|\mathcal{G}](\omega)],$ (A.8)
where the first and second equations are according to Step 1 and the third one is due to the definition of conditional expectation. Since both $\tilde{g}$ and $\mathbb{E}[h(\Lambda(\cdot),\cdot)|\mathcal{G}]$ are measurable with respect to $\mathcal{G}:=\sigma(\Lambda)\lor\mathcal{A}$, the arbitrary of $G$ implies
$g(\Lambda(\omega),\omega)=\mathbb{E}[h(\Lambda(\cdot),\cdot)|\mathcal{G}](\omega)\ \ a.s..$
Q.E.D
A.2 Proof of Results in Section 2
Proof of Theorem 1 We only focus on proving right tail of (8).
Step 1(Construction of super-martingale) Some simple algebra shows that, when $\lambda\in[0,\frac{1}{b}]$,
$e^{\lambda Z_{i}-\frac{\lambda^{2}}{2}Z_{i}^{2}}\leq 1+\lambda Z_{i}$ (A.9)
holds for any given $i$ and every $\omega\in\Omega$. Since $(Z_{i},\mathcal{A}_{i})_{i\geq 1}$ is a MDS, (A.9) implies
$\mathbb{E}\Big[\exp\Big(\lambda Z_{i}-\frac{\lambda^{2}Z_{i}^{2}}{2}\Big)|\mathcal{A}_{i-1}\Big]\leq 1+\lambda\mathbb{E}[Z_{i}|\mathcal{A}_{i-1}]=1\ \$(A.10)
holds almost surely for every $\lambda\in[0,1/b]$. For this arbitrary given $\lambda\in[0,1/b]$, we also define
$L_{m,\lambda}:=\exp\Big(\lambda\mathbf{M}_{m}-\frac{\lambda^{2}}{2}[\mathbf{M}]_{m}\Big).$ (A.11)
Then, using (A.10) and basic properties of MDS yields
$\mathbb{E}[L_{m,\lambda}|\mathcal{A}_{m-1}]=\mathbb{E}\Big[\exp\Big(\lambda\mathbf{M}_{m}-\frac{\lambda^{2}}{2}[\mathbf{M}]_{m}\Big)\Big|\mathcal{A}_{m-1}\Big]$
$=$ $\exp\Big(\lambda\mathbf{M}_{m-1}-\frac{\lambda^{2}}{2}[\mathbf{M}]_{m-1}\Big)\mathbb{E}\Big[\exp\Big(\lambda W_{m}-\frac{\lambda^{2}}{2}W_{m}^{2}\Big)\Big|\mathcal{A}_{m-1}\Big]$
$\leq$ $\exp\Big(\lambda\mathbf{M}_{m-1}-\frac{\lambda^{2}}{2}[\mathbf{M}]_{m-1}\Big)=L_{m-1,\lambda}.$
Thus, we manage to prove that, for any given $\lambda\in[0,1/b]$, $\{L_{m,\lambda},\mathcal{A}_{m}\}_{m\geq 1}$ is a super martingale such that
$\mathbb{E}[L_{m,\lambda}]\leq E[L_{1,\lambda}]\leq 1,\ \forall\ m\geq 1.$ (A.12)
Step 2 (Extension of filter and randomization of $\lambda$) By defining
$L_{m}(\lambda,\cdot):=\exp\Big(\lambda\mathbf{M}_{m}(\cdot)-\frac{\lambda^{2}}{2}[\mathbf{M}]_{m}(\cdot)\Big)\ \ \text{and}\ \ g_{m}(\lambda,\omega):=\mathbb{E}[L_{m}(\lambda,\cdot)](\omega),$
we have
  • (i)
    For each $m\geq 1$, both $g_{m}(\lambda,\omega)$ and $L_{m}(\lambda,\cdot)$ are real-valued and non-negative measurable mappings defined on $0,1/b]\times\Omega$;
  • (ii)
    $\max_{m\geq 1}\sup_{\lambda\in[0,1/b]}g_{m}(\lambda,\omega)\leq 1\$holds almost surely.
Here point (ii) is due to (A.12). Then, for any random variable $\Lambda$ taking value in $[0,1/b]$, by denoting $\mathcal{G}_{m}:=\sigma(\Lambda)\lor\mathcal{A}_{m}$ and $L_{m,\Lambda}:=L_{m}(\Lambda(\cdot),\cdot)$, $\{L_{m,\Lambda}\}_{m\geq 1}$ is adapted to filter $\{\mathcal{G}_{m}\}_{m\geq 0}$. $\mathcal{G}_{0}$ is any sigma algebra contained by the intersection of $\bigcap_{i\geq 1}\mathcal{G}_{t}$ and independent of $\sigma(\bigcup_{i\geq 1}\mathcal{G}_{i})$. This can always be done by properly enlarging the original probability space, like introducing product space. Meanwhile, Lemma A.4 implies that
$\mathbb{E}[L_{m,\Lambda}|\mathcal{G}_{m-1}](\omega)=\mathbb{E}[L_{m}(\Lambda(\cdot),\cdot)|\mathcal{G}_{m-1}](\omega)=g_{m}(\Lambda(\omega),\omega)\ \ \text{a.s.}$ (A.13)
holds for any given $t\geq 1$. Combining law of iterated expectation and point (ii) above yields
$\mathbb{E}[L_{m,\Lambda}]=\int_{\Omega}\mathbb{E}[L_{m,\Lambda}|\mathcal{G}_{m-1}](\omega)d\mathbb{P}=\int_{\Omega}g_{m}(\Omega(\omega),\omega)d\mathbb{P}\leq\max_{m\geq 1}\sup_{\lambda\in[0,1/b]}g_{m}(\lambda,\omega)\mathbb{P}(\Omega)\leq 1.$ (A.14)
By letting $m=n$, using with Markov inequality yields
$\mathbb{P}\Big(\Lambda\mathbf{M}_{n}-\frac{\Lambda^{2}}{2}[\mathbf{M}]_{n}\geq t\Big)=\mathbb{P}\Big(\exp\Big(\Lambda\mathbf{M}_{n}-\frac{\Lambda^{2}}{2}[\mathbf{M}]_{n}\Big)\geq e^{t}\Big)=\mathbb{P}(L_{n,\Lambda}\geq e^{t})\leq e^{-t}$ (A.15)
holds for any $t>0$, $n\geq 1$ and any random variable $\Lambda$ taking value in $[0,1/b]$. Step 3 Define
$\mathcal{E}_{t}=\Big\{\Lambda\mathbf{M}_{n}-\frac{\Lambda^{2}}{2}[\mathbf{M}]_{n}\geq t\Big\},\ \ E_{t}=\Big\{\mathbf{M}_{n}\geq\sqrt{2t([\mathbf{M}]_{n}\lor 2b^{2}t)}\Big\},\ \Lambda=\min\Big\{\sqrt{\frac{2t}{[\mathbf{M}]_{n}}},\frac{1}{b}\Big\}.$
Obviously, we can decompose event $E_{t}$ as
$E_{t}=E_{t}^{1}\cup E^{2}_{t},$
where $E_{t}^{1}=E_{t}\cap\{[\mathbf{M}]_{n}\geq 2b^{2}t\}$ and $E_{t}^{2}=E_{t}\cap\{[\mathbf{M}]_{n}<2b^{2}t\}$.
When $\omega\in E^{1}_{t}$, $\Lambda(\omega)=\sqrt{\frac{2t}{[\mathbf{M}]_{n}(\omega)}}$, $\mathbf{M}_{n}(\omega)\geq\sqrt{2[\mathbf{M}]_{n}(\omega)t}$ and $\mathbf{M}]_{n}(\omega)\geq 2b^{2}t$ hold simultaneously. Some basic algebra shows
$\Lambda(\omega)\mathbf{M}_{n}(\omega)-\frac{\Lambda^{2}(\omega)}{2}[\mathbf{M}]_{n}(\omega)\geq\Lambda(\omega)\sqrt{2[\mathbf{M}]_{n}(\omega)t}-\frac{\Lambda^{2}(\omega)}{2}[\mathbf{M}]_{n}(\omega)=2t-t=t,$
where the second last equality is based on the fact that $\Lambda(\omega)=\sqrt{\frac{2t}{[\mathbf{M}]_{n}(\omega)}}$. Thus, for any $t>0$,
$E^{1}_{t}\subset\mathcal{E}_{t}.$ (A.16)
When $\omega\in E^{2}_{t}$, since $\Lambda(\omega)=\frac{1}{b}$, $\mathbf{M}_{n}(\omega)\geq 2bt$ and $\mathbf{M}]_{n}(\omega)<2b^{2}t$ hold simultaneously and
$\Lambda(\omega)\mathbf{M}_{n}(\omega)-\frac{\Lambda^{2}(\omega)}{2}[\mathbf{M}]_{n}(\omega)\geq\frac{1}{b}2bt-\frac{1}{2b^{2}}2b^{2}t=2t-t=t,$
we obtain
$E^{2}_{t}\subset\mathcal{E}_{t}.$ (A.17)
Above all, combining (A.16) and (A.17) yields
$\mathbb{P}(E_{t})\leq\mathbb{P}(\mathcal{E}_{t})\leq e^{-t},\ \forall\ t>0.$
Q.E.D
Proof of Corollary 1 Since $(Z_{i})_{i\geq 0}$ is adapted to $(\mathcal{A}_{i})_{i\geq 0}$ and $\mathbb{E}[Z_{i}|\mathcal{A}_{i-1}]=\mu$, $(Z_{i}-\mu,\mathcal{A}_{i})_{i\geq 0}$ is a martingale difference sequence. Then, applying Theorem 1 indicates that, for any given $\delta>0$ and $n\geq 1$, the following event happens with probability no less than $1-2\delta$,
$\left\{\Big|\overline{\mathbf{M}}_{n}-\mu\Big|\leq\sqrt{\frac{2\log(\frac{1}{\delta})\sum_{i=1}^{n}(Z_{i}-\mu)^{2}}{n^{2}}}\lor\frac{2b_{V}\log(\frac{1}{\delta})}{n}\right\}=:E.$ (A.18)
Recall that, for each $n\geq 1$, we have $\sum_{i=1}^{n}(Z_{i}-\mu)^{2}=\sum_{i=1}^{n}(Z_{i}-\overline{\mathbf{M}}_{n})^{2}+n(\overline{\mathbf{M}}_{n}-\mu)^{2}$, and $\sqrt{x+y}\leq\sqrt{x}+\sqrt{y}$ hold for any $x,y\geq 0$. Some simple algebra shows that
$\sqrt{\frac{2\log(\frac{1}{\delta})\sum_{i=1}^{n}(Z_{i}-\mu)^{2}}{n^{2}}}\leq\sqrt{\frac{2\log(\frac{1}{\delta})\sum_{i=1}^{n}(Z_{i}-\overline{\mathbf{M}}_{n})^{2}}{n^{2}}}+\sqrt{\frac{2\log(\frac{1}{\delta})}{n}}|\overline{\mathbf{M}}_{n}-\mu|.$ (A.19)
Please note that, for any given $a,b,c,d\geq 0$ and $d\leq a$, event $\{a\leq\max\{b+d,c\}\}\subset\{a-d\leq b\lor c\}$. Thus,
$E_{\delta}\subset\left\{\Big(1-\sqrt{2\frac{1}{n}\log(\frac{1}{\delta})}\Big)\Big|\overline{\mathbf{M}}_{n}-\mu\Big|\leq\sqrt{\frac{2\log(\frac{1}{\delta})\sum_{i=1}^{n}(Z_{i}-\overline{\mathbf{M}}_{n})^{2}}{n^{2}}}\lor\frac{2b_{V}\log(\frac{1}{\delta})}{n}\right\}=E^{\prime}.$ (A.20)
Thus, we obtain $\mathbb{P}(E^{\prime})\geq P(E)\geq 1-2\delta$, which finish the proof.
Proof of Corollary 2 Since the proof of Corollary 2 is a direct applications of Theorem 1 and Corollary 2, we only show the key points here. Since $(Z_{i}-\mu,\mathcal{A}_{i})_{i\geq 0}$ is still a martingale difference sequence, According to (A.18), we construct event
$\mathcal{E}=\left\{\sqrt{\frac{2\log(\frac{1}{\delta}\sum_{i=1}^{n}(Z_{i}-\mu)^{2})}{n^{2}}}<\frac{2b_{V}\log(\frac{1}{\delta})}{n}\right\}.$ (A.21)
Since
$\sqrt{\frac{2\log(\frac{1}{\delta})\sum_{i=1}^{n}(Z_{i}-\mu)^{2}}{n^{2}}}=\sqrt{\frac{2\log(\frac{1}{\delta})[\frac{1}{n}\sum_{i=1}^{n}((Z_{i}-\mu)^{2}-\text{var}(Z_{i}))]}{n}+\frac{2\log(\frac{1}{\delta})\text{Var}(Z_{0})}{n}},$
some simple algebra shows that
$\mathcal{E}=\left\{\frac{1}{n}\sum_{i=1}^{n}((Z_{i}-\mu)^{2}-\text{var}(Z_{i}))+\text{Var}(Z_{0})<\frac{2\log(\frac{1}{\delta})b_{V}^{2}}{n}\right\}.$
For any given $\delta>0$ and $\eta\in(0,1)$, $N(\delta,\eta)=\min\{n\geq 1:\eta\text{Var}(Z_{0})\geq\frac{2\log(\frac{1}{\delta})b_{V}^{2}}{n}\}$. Then, for any $n\geq N(\delta,\eta)$, we have
$\mathcal{E}\subset\left\{\frac{1}{n}\sum_{i=1}^{n}(\text{Var}(Z_{i})-(Z_{i}-\mu)^{2})\geq(1-\eta)\text{Var}(Z_{0})\right\}=:\mathcal{E}^{\prime}.$
Above all, (A.18) implies
$\mathbb{P}(E)\leq\mathbb{P}(E\cap\mathcal{E}^{c})+\mathbb{P}(\mathcal{E})\leq\mathbb{P}(E\cap\mathcal{E}^{c})+\mathbb{P}(\mathcal{E}^{\prime})$
$\Rightarrow$ $\mathbb{P}(E\cap\mathcal{E}^{c})\geq\mathbb{P}(E)-\mathbb{P}(\mathcal{E}^{\prime})\geq 1-2\delta-\mathbb{P}(\mathcal{E}^{\prime})$ (A.22)
Since $\mathbb{E}[(Z_{i}-\mu)^{2}|\mathcal{A}_{i-1}]=\sigma^{2}>0$, using Freedman inequality yields
$\mathbb{P}(\mathcal{E}^{\prime})\leq\exp\left(-\frac{n(1-\eta)^{2}\sigma^{4}}{2(m_{4}+\frac{1}{3}b(1-\eta)\sigma^{2})}\right).$ (A.23)
Finally, combining (A.2), (A.23) and the calculation of (A.19) finishes the proof.
Proof of Theorems 2 and 3 The proof of Theorems 2 and 3 are nearly the same. So we give a proof which incorporates both cases. For convenience, we introduce notation
$\triangle_{n}$ $=:2(b_{n}-a_{n}),\ \Psi_{n}=:\Phi_{n},\ \$$\text{when we focus on $\phi$-mixing condition};$
$\triangle_{n}$ $=:B_{n},\ \Psi_{n}=:\widetilde{\Phi}_{n},$ $\text{when we focus on $\tilde{\phi}$-mixing condition}.$
Step 1 (Auxiliary triangular array and Gordin-coboundary decomposition)
For each given $n\geq 1$, introduce triangular array $\mathbf{Z}_{n}=\{Z_{nj}\}_{j\in\mathbb{Z}}$ such that
  • $(Z_{n1},...,Z_{nn})$ is identically distributed as $(Z_{1},...,Z_{n})$.
  • $\sigma(Z_{nk}:k\leq 0)$, $\sigma(Z_{n1},...,Z_{nn})$ and $\sigma(Z_{nk}:k\geq n+1)$ are mutually independent.
  • $\{Z_{nk}\}_{k\geq n+1}$ are copies of $Z_{nn}$.
According to ①, for any given $n\geq 1$, identity
$\mathbb{P}(T(Z_{n1},...,Z_{nn})\in A)=\mathbb{P}(T(Z_{1},...,Z_{n})\in A)$
holds for all measurable mapping $T:\mathcal{Z}^{n}\to\mathbb{R}$ and Borel set $A\subset\mathbb{R}$. Thus, we only need to focus on investigating the concentration phenomenon of $\sum_{i=1}^{n}(h_{n}(Z_{ni})-\mu_{n})=:S_{nn}(h_{n})$.
A noteworthy fact is that, by denoting $\mathcal{F}_{ni}=\sigma(Z_{nj}:j\leq i)$, ② and ③ above implies the following identity holds for every given $n\geq 1$,
$\max_{i\in\mathbb{Z}}\sum_{k>i}||\mathbb{E}[(h_{n}(Z_{nk})-\mu_{n})|\mathcal{F}_{ni}]||_{p}=\max_{i\in\mathbb{Z}}\sum_{k=i+1}^{n}||(\mathbb{E}[h_{n}(Z_{nk})-\mu_{n})|\mathcal{F}_{ni}]||_{p}<\infty.$ (A.24)
Thus, for each given $n\geq 1$, Lemma A.6 implies the following decomposition
$h_{n}(Z_{ni})$ $=D_{ni}(h_{n})-\varepsilon_{ni}(h_{n})+\varepsilon_{n,i-1}(h_{n}),$ (A.25)
$S_{nN}(h_{n})$ $=\sum_{i=1}^{N}D_{ni}(h_{n})-\varepsilon_{nN}(h_{n})+\varepsilon_{n0}(h_{n})=\sum_{i=1}^{N}D_{ni}(h_{n}),\ \ \forall\ N\geq 1.$ (A.26)
where
$D_{ni}(h_{n})$ $=\sum_{k=i}^{n}(\mathbb{E}[h_{n}(Z_{nk})|\mathcal{F}_{ni}]-\mathbb{E}[h_{n}(Z_{nk})|\mathcal{F}_{n,i-1}])=\sum_{k=1}^{n}(\mathbb{E}[h_{n}(Z_{nk})|\mathcal{F}_{ni}]-\mathbb{E}[h_{n}(Z_{nk})|\mathcal{F}_{n,i-1}]),$
$\varepsilon_{ni}(h_{n})$ $=\sum_{k>i}\mathbb{E}[(h_{n}(Z_{nk})-\mu_{n})|\mathcal{F}_{ni}],\ \ \varepsilon_{nn}(h_{n})=\varepsilon_{n0}(h_{n})=0.$
The second inequality (A.26) is because, according to ②, $\varepsilon_{nn}(h_{n})=\varepsilon_{n0}=0$ holds almost surely.
Additionally, for each $i$ and $n$, we have
[…]
Step 2 (Block) Based on the $m$, $l$ and $B_{j}$’s introduced in Theorem 2, by defining $a_{j}=(j-1)[l]+1$ and $b_{j}=a_{j}+[l]-1$ for $1\leq j\leq m-1$, we have
$B_{j}=\left\{\begin{aligned} &\{a_{j}+1,...,b_{j}\},\ \ &1\leq j\leq m;\\ &\{m[l]+1,...,n\},&j=m+1.\end{aligned}\right$. (A.28)
Hence, together with (A.26), we obtain
$S_{nb_{m}}(h_{n})=\sum_{i=1}^{b_{m}}D_{ni}(h_{n})=\sum_{j=1}^{m}(\sum_{i\in B_{j}}D_{ni}(h_{n}))=:\sum_{j=1}^{m}M_{nj},$ (A.29)
which yields
$S_{nn}(h_{n})=S_{nb_{m}}(h_{n})+\sum_{i=m[l]+1}^{n}(h_{n}(Z_{ni})-\mu_{n})=:S_{nb_{m}}(h_{n})+r_{n}.$ (A.30)
Meanwhile, (A.25) implies that
$M_{nj}=\sum_{i\in B_{j}}D_{ni}(h_{n})=\left\{\begin{aligned} &\sum_{i\in B_{j}}(h_{n}(Z_{ni})-\mu_{n})+\varepsilon_{nb_{j}}(h_{n})-\varepsilon_{n,a_{j}-1}(h_{n}),\ \ &2\leq j\leq m;\\ &\sum_{i=m[l]+1}^{n}(h_{n}(Z_{ni})-\mu_{n})-\varepsilon_{n,m[l]}(h_{n}),&j=m+1;\\ &\sum_{i=1}^{[l]}(h_{n}(Z_{ni})-\mu_{n})+\varepsilon_{n,[l]}(h_{n}),&j=1.\end{aligned}\right$. (A.31)
Define $\mathcal{G}_{nj}:=\mathcal{F}_{nb_{j}}$, for $0\leq j\leq m$, and $G_{nm}=\mathcal{F}_{nn}$, for $j=m+1$. Thus, $\{M_{nj}\}_{j=0}^{m}$ is adapted to $\{\mathcal{G}_{nj}\}_{j=0}^{m}$. Furthermore, some simple algebra shows
$\mathbb{E}[M_{nj}|\mathcal{G}_{n,j-1}]=\mathbb{E}[M_{nj}|\mathcal{F}_{n,a_{j}-1}]=\sum_{i=a_{j}}^{b_{j}}\mathbb{E}[D_{ni}(h_{n})|\mathcal{F}_{n,a_{j}-1}]$
$=$ $\sum_{k=1}^{n}\mathbb{E}[\sum_{i=a_{j}}^{b_{j}}(\mathbb{E}[h_{n}(Z_{nk})|\mathcal{F}_{ni}]-\mathbb{E}[h_{n}(Z_{nk})|\mathcal{F}_{n,i-1}])|\mathcal{F}_{n,a_{j}-1}]$
$=$ $\sum_{k=1}^{n}\mathbb{E}[(\mathbb{E}[h_{n}(Z_{nk})|\mathcal{F}_{nb_{j}}]-\mathbb{E}[h_{n}(Z_{nk})|\mathcal{F}_{n,a_{j}-1}])|\mathcal{F}_{n,a_{j}-1}]$
$=$ $\sum_{k=1}^{n}(\mathbb{E}[h_{n}(Z_{nk})|\mathcal{F}_{n,a_{j}-1}]-\mathbb{E}[h_{n}(Z_{nk})|\mathcal{F}_{n,a_{j}-1}])=0,$
where the last equality is due to tower property of conditional expectation. Therefore, $\{(M_{nj},\mathcal{G}_{nj})\}_{j=1}^{m}$ is martingale difference triangular array. Together with (A.30), we manage to write the partial sum of mixing process as partial sum of martingale difference plus a remainder.
(Step 3) (Moment Conditions of $M_{nj}$) This step is dedicated to calculate some important moment conditions of $M_{nj}$ so that we can easily employ Theorem 1 in the next step.
(Upper bound of $M_{nj}||_{\infty}$) Some basic algebra shows that, under both $\phi$ and $\tilde{\phi}-$mixing conditions, we have
$|M_{nj}|\leq\max_{1\leq j\leq m}\Big|\sum_{i\in B_{j}}(h_{n}(Z_{ni})-\mu_{n})+\varepsilon_{nb_{j}}(h_{n})-\varepsilon_{n,a_{j}-1}(h_{n})\Big|\leq[l](b_{n}-a_{n})+\triangle_{n}\Psi_{n}.$ (A.32)
(Upper bound of $\sqrt{\sum_{j=1}^{m}M_{nj}^{2}}$) By denoting $T_{nj}=\sum_{i\in B_{j}}(h_{n}(Z_{ni})-\mu_{n})$ and $R_{nj}=\varepsilon_{nb_{j}}(h_{n})-\varepsilon_{n,a_{i}-1}(h_{n})$,
$\sum_{j=1}^{m}M^{2}_{nj}=\sum_{j=1}^{m}(T_{nj}+R_{nj})^{2}$
$\leq$ $\sum_{j=1}^{m}(|T_{nj}|^{2}+|R_{nj}|^{2}+2|T_{nj}R_{nj}|)$
$\leq$ $(1+\eta)(\sum_{j=1}^{m}|R_{nj}|^{2}+\xi_{n})+(1+\frac{1}{\eta})(\sum_{j=1}^{m}|T_{nj}|^{2})$
holds for any given $\eta>0$ and arbitrary $\xi_{n}\searrow 0$. A noteworthy fact is that, for any given $A\geq 0$ and $B>0$, $f(\eta)=(1+\eta)A+(1+\frac{1}{\eta})B$ obtains its minimum value at point $\eta=\frac{B}{A}$, which asserts
$f_{\min}=f(\frac{B}{A})=2\sqrt{AB}+A+B=(\sqrt{A}+\sqrt{B})^{2}.$
Since $\sum_{j=1}^{m}|R_{nj}|^{2}+\xi_{n}>0$ holds for every $n$, by regarding it as the term $A$ above, we can show that
$\sqrt{\sum_{j=1}^{m}M_{nj}^{2}}\leq\sqrt{\sum_{j=1}^{m}T_{nj}^{2}}+\sqrt{\sum_{j=1}^{m}R_{nj}^{2}+\xi_{n}}\leq\sqrt{\sum_{j=1}^{m}T_{nj}^{2}}+\sqrt{\sum_{j=1}^{m}R_{nj}^{2}}+\sqrt{\xi_{n}}$ (A.33)
holds almost surely for arbitrary $\xi_{n}\searrow 0$, where the last inequality is due the fact that $\sqrt{x+y}\leq\sqrt{x}+\sqrt{y}$ holds for every $x,y\geq 0$. Note that (A.27) implies
$\sqrt{\sum_{j=1}^{m}R_{nj}^{2}}\leq 2\triangle_{n}\Psi_{n}\sqrt{m}.$ (A.34)
As for $\sum_{j=1}^{m}T_{nj}^{2}$, by letting $\bar{H}_{n}=\frac{1}{m[l]}\sum_{i=1}^{m{l}}h_{n}(Z_{ni})$, Lemma A.1 yields
$\sum_{j=1}^{m}T_{nj}^{2}$ $=\sum_{j=1}^{m}(\sum_{i\in B_{i}}(h_{n}(Z_{ni})-\bar{H}_{n}))^{2}+[l]^{2}m(\frac{1}{m[l]}S_{nb_{m}}(h_{n}))^{2}$
$=\sum_{j=1}^{m}(\sum_{i\in B_{i}}(h_{n}(Z_{ni})-\bar{H}_{n}))^{2}+\frac{n^{2}}{m}(\frac{1}{n}S_{nb_{m}}(h_{n}))^{2}$
which asserts
$\sqrt{\sum_{j=1}^{m}T_{nj}^{2}}\leq\sqrt{\sum_{j=1}^{m}(\sum_{i\in B_{i}}(h_{n}(Z_{ni})-\bar{H}_{n}))^{2}}+\Big|\frac{1}{n}S_{nb_{m}}(h_{n})\Big|\frac{n}{\sqrt{m}}.$ (A.35)
Combining (A.33)-(A.35), we obtain
$\sqrt{\sum_{j=1}^{m}M_{nj}^{2}}\leq\sqrt{\sum_{j=1}^{m}(\sum_{i\in B_{i}}(h_{n}(Z_{ni})-\bar{H}_{n}))^{2}}+\Big|\frac{1}{n}S_{nb_{m}}(h_{n})\Big|\frac{n}{\sqrt{m}}+2\triangle_{n}\Psi_{n}\sqrt{m}+\sqrt{\xi_{n}}.$ (A.36)
Step 4 (Application of Theorem 1) According to Theorem 1, we can show that, for any $\delta>0$ and arbitrary $\xi_{n}\searrow 0$, the following inequality holds with probability no less than $1-2\delta$,
$|S_{nb_{m}}(h_{n})|\leq\max\left\{\sqrt{2\sum_{j=1}^{m}M_{nj}^{2}\log(\frac{1}{\delta})},2\max_{1\leq j\leq m}|M_{nj}|\log(\frac{1}{\delta})\right\}\leq\max\{A,B\},$
where
$A=$ $\sqrt{2\log(\frac{1}{\delta})\sum_{j=1}^{m}(\sum_{i\in B_{i}}(h_{n}(Z_{ni})-\bar{H}_{n}))^{2}}+\Big|\frac{1}{n}S_{nb_{m}}(h_{n})\Big|n\sqrt{\frac{2\log(\frac{1}{\delta})}{m}}$
$+2\triangle_{n}\Psi_{n}\sqrt{2\log(\frac{1}{\delta})m}+\sqrt{2\log(\frac{1}{\delta})\xi_{n}},$
$B=$ $2([l](b_{n}-a_{n})+\triangle_{n}\Psi_{n}).$
Thus, some simple algebra yields
$\left(1-\sqrt{\frac{2\log(\frac{1}{\delta})}{[n/[l]]}}\right)\Big|\frac{1}{n}S_{nb_{m}}(h_{n})\Big|\leq\max\{A^{\prime},B^{\prime}\},$
where
$A^{\prime}$ $=\sqrt{2\log(\frac{1}{\delta})\frac{1}{n^{2}}\sum_{j=1}^{m}(\sum_{i\in B_{i}}(h_{n}(Z_{ni})-\bar{H}_{n}))^{2}}+2\triangle_{n}\Psi_{n}\sqrt{\frac{2\log(\frac{1}{\delta})m}{n^{2}}}+\sqrt{\frac{2\log(\frac{1}{\delta})\xi_{n}}{n^{2}}},$
$B^{\prime}$ $=\frac{2([l](b_{n}-a_{n})+\triangle_{n}\Psi_{n})}{n}.$
Meanwhile, regarding $|\frac{1}{n}r_{n}|\leq\frac{[l]}{n}(b_{n}-a_{n})$,
$\mathbb{P}\left(\Big|\frac{1}{n}S_{nn}(h_{n})\Big|\leq\tilde{\nu}_{n}(\delta)\max\{A^{\prime},B^{\prime}\}+\frac{[l]}{n}(b_{n}-a_{n})\right)\geq 1-2\delta,$
where $\tilde{\nu}_{n}(\delta)=\Big(1-\sqrt{\frac{2\log(\frac{1}{\delta})}{[n/[l]]}}\Big)^{-1}$.
(Proof of Theorems 4) Regarding the following proof is just modification of the proof of Theorems 2 and 3, we only highlight the key steps and omit the repetitive parts.
Step 1 (Crucial concentration phenomenon) Based on the $r_{n}=\sum_{i=m[l]+1}^{n}(h_{n}(Z_{ni})-\mu)$ introduced in (A.29) and the facts that $\mathbb{E}[r_{n}]=0$ and $\frac{r_{n}}{n-m[l]}\in[a_{n},b_{n}]$, for any $\tau_{1n}\nearrow\infty$ and $t>0$, we have
$\mathbb{P}\Big(\Big|\frac{1}{\tau_{1n}}r_{n}\Big|>t\Big)=\mathbb{P}\Big(\Big|\frac{1}{n-m[l]}r_{n}\Big|>\frac{\tau_{1n}}{(n-m[l])}t\Big)\leq 2\exp\Big(-2\Big(\frac{(\tau_{1n}/[l])t}{(b_{n}-a_{n})}\Big)^{2}\Big),$ (A.37)
where $C_{n}=\frac{n}{n-m[l]}\geq\frac{n}{[l]}$. Similarly, according to the definition of $T_{nj}$ and $R_{nj}$ defined in Step 3 of the proof of Theorems 2 and 3, we have $\mathbb{E}[T_{nj}]=\mathbb{E}[R_{nj}]=0$ and $|T_{nj}|\leq[l](b_{n}-a_{n})$, $|R_{nj}|\leq 2\triangle_{n}\Psi_{n}$. Thus, for any given $\tau_{2n}\nearrow\infty$ and $t>0$, we have
$\mathbb{P}\Big(\max_{1\leq j\leq m}\frac{1}{\tau_{2n}}|R_{nj}|>t\Big)$ $\leq 2m\exp\left(-0.5\Big(\frac{\tau_{2n}t}{\triangle_{n}\Psi_{n}}\Big)^{2}\right).$ (A.38)
Step 2 (Generic high-probability bound) According to the proof of Theorems 2 and 3, we have generic high-probability (probability larger than $1-2\delta$) bound as follow,
$|S_{nn}(h_{n})|\leq|S_{nb_{m}}(h_{n})|+|r_{n}|=|\sum_{j=1}^{m}M_{nj}|+|r_{n}|$
$\leq$ $\max\left\{\sqrt{2\log(\frac{1}{\delta})}\left(\sqrt{\sum_{j=1}^{m}T^{2}_{nj}}+\sqrt{\sum_{j=1}^{m}R^{2}_{nj}}+\sqrt{\xi_{n}}\right),2\log(\frac{1}{\delta})(\max_{1\leq j\leq m}|T_{nj}|+\max_{1\leq j\leq m}|R_{nj}|)\right\}+|r_{n}|.$
Then, some simple algebra shows that, for any given $\delta>0$ and arbitrary $\xi_{n}\searrow 0$, we have
$\mathbb{P}\Big(|\frac{1}{n}S_{nn}(h_{n})|\leq\tilde{\nu}_{n}(\delta)\max\{\mathbf{A},\mathbf{B}\}+|\frac{r_{n}}{n}|\Big)\geq 1-2\delta,$ (A.39)
$\mathbf{A}=$ $\sqrt{\frac{2\log(\frac{1}{\delta})[\sum_{j=1}^{m}(\sum_{i\in B_{j}}h_{n}(Z_{ni})-\bar{H}_{n})^{2}]}{n^{2}}}+\frac{\max_{1\leq j\leq m}|R_{nj}|}{\sqrt{n[l]}}\sqrt{2\log(\frac{1}{\delta})}+\sqrt{\frac{2\xi_{n}\log(\frac{1}{\delta})}{n^{2}}},$
$\mathbf{B}=\frac{2\log(\frac{1}{\delta})[l](b_{n}-a_{n})}{n}+\frac{2\log(\frac{1}{\delta})}{n}\max_{1\leq j\leq m}|R_{nj}|.$
Note that (A.37) implies, for any user-defined $c_{n}\searrow 0$,
$\mathbb{P}\Big(\Big|\frac{1}{n}r_{n}\Big|>c_{n}\Big)\leq 2\exp\left(-2\Big(\frac{(n/[l])c_{n}}{(b_{n}-a_{n})}\Big)^{2}\right).$ (A.40)
(A.38) indicates that, for any user-defined $t_{n}\searrow 0$ and $\theta^{\prime}\in(0,\frac{1}{2})$,
$\mathbb{P}\Big(\max_{1\leq j\leq m}\frac{|R_{nj}|}{\sqrt{n[l]}}>t_{n}\Big)$ $\leq 2m\exp\left(-0.5\Big(\frac{\sqrt{n[l]}t_{n}}{\triangle_{n}\Psi_{n}}\Big)^{2}\right);$ (A.41)
$\mathbb{P}\Big(\max_{1\leq j\leq m}\frac{|R_{nj}|}{n^{\theta^{\prime}}}>\frac{1}{2}\Big)$ $\leq 2m\exp\left(-\Big(\frac{n^{\theta^{\prime}}}{16\triangle_{n}\Psi_{n}}\Big)^{2}\right).$ (A.42)
Combining (A.39)-(A.42), for user-defined $\theta\in(0,\frac{1}{6}]$, $\theta^{\prime}\in(0,\frac{1}{2})$ and $t_{n}\searrow 0$, we have
$\mathbb{P}\Big(|\frac{1}{n}S_{nn}(h_{n})|\leq\tilde{\nu}_{n}(\delta)\max\{\mathbf{A}^{\prime},\mathbf{B}^{\prime}\}+\frac{1}{n^{\frac{1}{2}+\theta}}\Big)\geq 1-2\delta-\sum_{k=1}^{3}\text{Error}_{k},$ (A.43)
$\text{where}\ \mathbf{A}^{\prime}$ $=\sqrt{\frac{2\log(\frac{1}{\delta})[\sum_{j=1}^{m}(\sum_{i\in B_{j}}h_{n}(Z_{ni})-\bar{H}_{n})^{2}]}{n^{2}}}+t_{n}\sqrt{2\log(\frac{1}{\delta})}+\sqrt{\frac{2\xi_{n}\log(\frac{1}{\delta})}{n^{2}}},$
$\mathbf{B}^{\prime}$ $=\frac{2\log(\frac{1}{\delta})[l](b_{n}-a_{n})}{n}+\frac{\log(\frac{1}{\delta})}{n^{1-\theta^{\prime}}},$
and $\text{Error}_{k}$, $k=1,2,3$, coincides with the upper bounds in (A.40)-(A.42).